OpenStudy (helpmeplease14):

A circular swimming pool holds 1.22*10^6 cubic inches of water. It is being filled at a rate of 1.5*10^3 cubic inches per minute. How many hours will it take to fill the swimming pool?

OpenStudy (astrophysics):

Since you're given the rate per minute, you can divide 1.22*10^6 by 1.4*10^3 to see how many minutes it will take then simply divide that by 60 to put it in hours

OpenStudy (helpmeplease14):

K

OpenStudy (anonymous):

no. of minutes to fill the pool\[=\frac{ 1.22*10^6 }{ 1.5*10^3 }\]

OpenStudy (astrophysics):

1.5* not 1.4

OpenStudy (helpmeplease14):

Yeah ik

OpenStudy (helpmeplease14):

8.13 × 10^4?

OpenStudy (astrophysics):

\[\frac{ 1.22 \times 10^6 }{ 1.5 \times 10^3} \times 1 \min \times \frac{ 1 hr }{ 60 \min }\] we have that 1 minute because it's 1.5*103 in^3/min so then we use our algebra rules to flip and multiply

OpenStudy (astrophysics):

How did you get 10^4?

OpenStudy (helpmeplease14):

Because I settee up the factors and the powers so it's (1.22/ 1.5) (10^6/ 10^3) Divided the factors and got 0.813 0.813 (10^6-^3) I subtracted the exponents and got 3 0.813 × 10^3 The answer has to be in scientific notation so it's 8.13 × 10^4 and added the exponent by 1 to get 4

OpenStudy (astrophysics):

No it would become 8.13*10^2 then it does not increase

OpenStudy (helpmeplease14):

Oh

OpenStudy (astrophysics):

8.13*10^4 = 81300

OpenStudy (astrophysics):

0.813*10^3 = 813 = 8.13*10^2

OpenStudy (astrophysics):

But you want it in hours still, so make sure you divide by 60

OpenStudy (helpmeplease14):

So 813/60

OpenStudy (astrophysics):

Yup

OpenStudy (helpmeplease14):

13.55 or 13 5/9

OpenStudy (astrophysics):

Lol it's the same thing, but what ever it's asking for

OpenStudy (astrophysics):

Wait no 13.55, 13.55... is 13 5/9

OpenStudy (astrophysics):

5/9 would be if the 5555555 was repeating

OpenStudy (helpmeplease14):

I know they're the same I'm just showing in decimal form and fraction form

OpenStudy (helpmeplease14):

Ty for the help

OpenStudy (astrophysics):

Yw